Calculus I Fall 2001 MA113:13-15, 19-21 Russell Brown

Below are the first few homework assignments from Chapter 3.

• Homework #9. Due Wednesday, 24 October 2001. §3.1 #12, 24. §3.2 #8, 18, 20.

• Homework D. Due Friday, 26 October 2001. §3.1 #68. In part a), please give conditions on the coefficients a, b, c and d which can be used to determine how many critical numbers a cubic polynomial has. (Hint: Discriminant.) These conditions can be used to help find examples.

• Homework #10. Due Monday, 29 October 2001. §3.3 #8, 18, 36, 38. §3.4 #8, 24, 18.

• Homework E. Due Wednesday, 31 October 2001. §3.3 #32.

October 16, 2001